A solid sphere of radius 10.0 cm is made of copper. What is its moment of inertia about an axis through the center of mass? The density of copper is 8.89 × 103 kg/m3 .


1
Expert's answer
2020-09-21T08:29:03-0400

"I=\frac{2}{5}mR^2=\frac{2}{5}R^2\cdot \frac{4}{3}\pi\rho R^3=\frac{8}{15}\pi\rho R^5"

"I=\frac{8}{15}\pi(8890)(0.1)^5=0.149\ kg\cdot m^2"